translate this word problem as a system of equations and then solve using substitution.: find two numbers whose sum is 58 and whose difference is 16.

Answer by checkley79(3341)

X+Y=58
X-Y=16 ADD.
--------------------
2X=74
X=74/2
X=37 ANS. FOR THE LARGER NUMBER.
37+Y=58
Y=58-37
Y=21 ANS. FOR THE SMALLER NUMBER.
PROOF:
37-21=16
16=16